OpenStudy (anonymous):

Can someone explain to me how to factor this: y(3y+11)=20

OpenStudy (lalaly):

\[3y^2 +11y-20 =0\]\[(y+5)(3y-4) =0\]

OpenStudy (anonymous):

muliply : 3y^2+11y -20 =0 ===> (3y -4 ) (y+5) = 0 ===> y=4/3 or y=-5